OpenStudy (anonymous):

[2.07] How many solutions does the equation 2|x − 3| + 4 = 4 have?

OpenStudy (anonymous):

lx-3l = 0 \[(x-3)^{2}=0\] x=3 Therfore only one solution

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

sure thing :) ask away!

OpenStudy (anonymous):

ok im trying to understand how to solve solutions

OpenStudy (anonymous):

for example |x − 4| = 6

OpenStudy (anonymous):

ok u see to get rid of thr modulus sign u have to square both sides of the equation

OpenStudy (anonymous):

so 4^2 = 6^2 ?

OpenStudy (anonymous):

\[(x-4)^{2}= 6^{2}\]